OpenStudy (anonymous):

evaluate lim(x->0) (sin^2x -x^2)/(x^2* sin^2x)

OpenStudy (anonymous):

yea way too early for this one

OpenStudy (amistre64):

this tends to involve trying to finagle a sin(x)/x into the picture

OpenStudy (anonymous):

if we sub x=0 the function = 0/0 indeterminate so i think we need to use hopital's rule

OpenStudy (anonymous):

i must admit i'm pretty rusty at limits

OpenStudy (amistre64):

\[\lim_{(x->0)} \frac{sin^2x -x^2}{x^2\ sin^2x}\] \[\lim_{(x->0)} \frac{(sin(x) -x)(sin(x) +x)}{x\ sin(x)x\ sin(x)}\] \[\lim_{(x->0)} \frac{sin(x) +x}{x\ sin(x)}*\frac{sin(x) -x}{x\ sin(x)}\] \[\lim_{(x->0)} \frac{sin(x)}{x\ sin(x)}+\frac{x}{x\ sin(x)}*\frac{sin(x)}{x\ sin(x)}-\frac{x}{x\ sin(x)}\] \[\lim_{(x->0)} \frac{1}{x}+\frac{1}{sin(x)}*\frac{1}{x}-\frac{1}{sin(x)}\] apparently i cant recall a nice way to do it this morning :)

OpenStudy (anonymous):

its not coming by l'hospital...

OpenStudy (amistre64):

ugh..wolfram maybe?

OpenStudy (anonymous):

http://www4d.wolframalpha.com/Calculate/MSP/MSP140119h4bf22d81d1h2100003ihaei7e167cf742?MSPStoreType=image/gif&s=59&w=488&h=2528 Pretty Long Solution .....Need a better Way........ *BookMark*